# Top Management Journal Portal: A Real-Source Search and Research Analytics Artifact for UTD-24 and FT50 Journals

### 09 Apr 2026 v1

@@ Abstract
This paper presents Top Management Journal Portal, a deployable web artifact for searching, monitoring, and interpreting literature from elite business and management journals. The system integrates the UTD-24 and Financial Times 50 (FT50) journal pools, retrieves live article metadata from the Crossref REST API, and organizes scholarly work into an end-to-end workflow spanning query formulation, result filtering, hotspot extraction, citation export, favorites management, and usage analytics. Unlike static journal directories or general-purpose academic search engines, the artifact is explicitly scoped to high-status management outlets and is designed to support sensemaking tasks that matter to researchers, doctoral students, and lab managers: identifying recent work, surfacing topical concentration, and converting search output into actionable research material. Architecturally, the system emphasizes source transparency, modularity, and low-cost public deployability through a lightweight Node.js service layer, a multi-page client interface, optional large-language-model enhancement for hotspot rewriting, and a free-tier persistence path through Supabase. The paper contributes both a functioning design artifact and an extensible architectural pattern for journal-pool-specific scholarly discovery, with implications for digital research infrastructure in information systems and business scholarship.
Keywords: design science, scholarly search, research analytics, business journals, Crossref, digital research infrastructure
@@ 1 Introduction
This paper introduces Top Management Journal Portal, a web-based research artifact designed to reduce that fragmentation. The artifact does not attempt to replace publisher websites or commercial bibliographic databases. Rather, it provides a focused, domain-specific decision support layer that combines journal-pool curation, real-source metadata retrieval, topic extraction, export utilities, and lightweight behavioral analytics. In design-science terms, the artifact addresses a class of recurrent scholarly search and sensemaking problems by creating a practical digital object for use in a specific problem domain (Hevner et al., 2004; Gregor and Hevner, 2013). The broader ambition is to show that even a relatively lightweight artifact can offer meaningful infrastructure value when its scope, sourcing logic, and task support are carefully aligned.
@@ 3 Design Objectives
Practical research support. Citation export, favorites, downloadable CSV output, and quick article-level commentary are included because they reduce friction in everyday research tasks.
@@ 4.1 Front-End Structure
    Article Search: Crossref retrieval, filters, hotspot analysis, export, favorites, and citation support.
@@
    Resources: conferences, submission tools, and external scholarly resources.
@@
    Analytics: visits, top search keywords, favorites ranking, trends, and traffic map widgets.
@@ 4.2 Back-End Services
    metadata retrieval and normalization,
@@
Presentation layer HTML, CSS, vanilla JavaScript Multi-page interface for search, journal browsing, export utilities, and analytics views
Search interaction Browser-side form state and query serialization Stable query submission, shareable URLs, local favorites, and citation actions
@@
Application layer Node.js HTTP server Static asset delivery, API routing, Crossref orchestration, and result normalization
Metadata source Crossref REST API Live retrieval of paper-level metadata, dates, authors, affiliations, and DOI records
@@
Presentation Layer Home    Journal Directory    Article Search    Resources    Analytics Application Layer Node.js server for routing, Crossref request orchestration, metadata normalization, query filtering, pagination, citation formatting, and analytics aggregation Analytics and Enrichment Layer Hotspot extraction    distributional analytics    citation export    favorites ranking    optional LLM enhancement Curated Sources UTD-24 metadata FT50 metadata External Metadata Crossref REST API Persistence Supabase analytics store or local fallback
Figure 1: System architecture of the Top Management Journal Portal artifact
@@ 5 Search-to-Insight Workflow
Given a keyword and optional filters, the back end first identifies the journals that belong to the selected pool. It then issues bounded Crossref requests journal by journal, normalizes the returned data into a common article schema, applies filtering and sorting logic, and computes analytical summaries from the resulting set.
@@
The implemented interface focuses on a small number of stable research tasks: choosing a journal pool, launching a live query, inspecting recent papers, extracting hotspots, and exporting references. Instead of emphasizing low-fidelity mockups, the report highlights the implemented views that researchers actually interact with during search and analysis. Figures 35 therefore serve not only as screenshots, but as evidence of how the artifact translates architectural intentions into concrete researcher-facing workflows.
@@
Figure 4: Implemented search interface showing live article retrieval, hotspot analysis, citation tools, and filtering controls
Figure 5: Implemented analytics interface with trend views, source tracking, favorites ranking, and traffic visualization
Figure 3 illustrates the artifact’s front-door logic: it foregrounds source transparency, journal-pool framing, and navigational shortcuts. Figure 4 shows the core operational surface, where live metadata retrieval, filtering, citation export, and hotspot analysis are tightly coupled in a single page. Figure 5 extends the artifact beyond search by exposing a monitoring layer for visits, usage concentration, and interaction patterns. Together, these implemented views demonstrate that the artifact is not only technically functional but also structurally aligned with recurring scholarly work practices.
@@ 6.1 Hotspot Extraction
Hotspot phrases are derived from titles and abstracts of the current result set. Candidate phrases are scored and then grouped into interpretable categories such as AI/Algorithms, Consumer/Market, Operations/Supply Chain, Finance/Accounting, and Strategy/Organization. When no native article keywords are available, the system falls back to rule-based extraction. When an LLM endpoint is configured, hotspot labels and quick takes can be rewritten into more readable summaries.
@@ 7.1 Technology Stack
The artifact uses Node.js for server-side routing and API orchestration, vanilla JavaScript for front-end logic, HTML/CSS for multi-page rendering, ECharts for result-set and analytics visualization, Supabase for free-tier persistence of usage analytics, and Render as a practical public deployment target. The implementation deliberately avoids a heavier front-end framework in order to minimize setup friction and hosting complexity. This restraint is important for the artifact’s design logic: a research support system is more likely to be adopted, replicated, and extended when the technical barrier to deployment remains low.
@@ 7.2 Front-End and Interaction Design
The front end is organized as a modular multi-page interface rather than a monolithic single-page application. This choice simplifies routing, page ownership, and public hosting while preserving a coherent user journey across home, journal browsing, live search, resource access, and analytics. Browser-side logic manages form state, URL synchronization, favorites persistence, citation actions, and optional user-provided large-language-model configuration. From a software-engineering perspective, this structure improves maintainability by keeping task-specific rendering concerns loosely coupled while still allowing shared interface patterns across pages.
@@ 7.3 Back-End Retrieval and Normalization
The back end mediates between curated journal pools and live metadata retrieval. For each query, the service first resolves the relevant journal set, then issues bounded Crossref requests on a journal-by-journal basis, normalizes returned records into a common schema, removes duplicates, applies filters, and computes analytical summaries. This orchestration layer is central to the artifact’s contribution: it operationalizes prestige-based journal curation as a search boundary, thereby aligning retrieval behavior with the evaluative logic commonly used in business schools.
@@ 7.4 Data Provenance and Source Integrity
The artifact is intentionally conservative about data provenance. Journal pools are anchored in documented UTD-24 and FT50 sources, while paper-level content is derived from Crossref records and exposed through DOI links and source labels. When article keywords are unavailable from native metadata, the system falls back to explicit heuristic extraction rather than presenting opaque or untraceable enrichment. Optional LLM support is treated as a presentation enhancement layer rather than a source of authoritative bibliographic truth.
@@ 7.6 Citation and Export Support
Each article object can be exported or cited through several output modes, including BibTeX, APA, MLA, Chicago, and plain text. CSV export and report export are also supported. These functions position the artifact closer to an everyday research tool than a demonstration-only prototype. In practice, they reduce a common break in scholarly workflow: the transition from discovering relevant papers to incorporating them into notes, bibliographies, and downstream review documents.
@@ 8 Discussion
The artifact contributes to design science and scholarly infrastructure in three ways. First, it demonstrates a domain-specific scholarly search artifact centered on elite business journals rather than undifferentiated bibliographic breadth. Second, it shows how real-source metadata and lightweight analytics can be combined into a usable search-to-insight workflow without requiring proprietary infrastructure. Third, it offers a deployable pattern for small research labs that need a manageable public-facing tool. More broadly, the artifact suggests that there is substantial room for middle-layer scholarly systems: systems that do not compete with large bibliographic platforms on scale, but outperform them on domain alignment, transparency, and task focus.
At the same time, the artifact has important limitations. Coverage depends on Crossref metadata availability, which varies across journals and articles. The hotspot layer is optimized for interpretability rather than formal topic-model validity. The system does not yet support collaborative annotation, full-text indexing, or citation-network exploration. Finally, the artifact has not yet undergone formal user evaluation with faculty or doctoral students. These limits should be interpreted as research opportunities rather than as disqualifying weaknesses: they point to the next design cycle needed to move from promising artifact to evaluated scholarly infrastructure.
@@ 10 Conclusion
Top Management Journal Portal is a focused research artifact for discovering, searching, and analyzing literature from elite business journals. By integrating UTD-24 and FT50 curation with live Crossref retrieval, lightweight analytical summaries, optional AI enhancement, citation utilities, and deployable analytics infrastructure, the artifact offers a practical response to the fragmented search workflows often faced by management researchers. Although it remains an early-stage artifact rather than a finished scholarly platform, it already demonstrates a credible and extensible design pattern for business-school-oriented digital research infrastructure. For ICIS-style design and artifact discussions, the promise of the system lies not only in what it currently does, but in the evaluative and infrastructural research program it makes possible.
@@ Acknowledgments
The current public-facing artifact credits Designed by Chuang Zhao (ADM Lab). The implementation also integrates external data and infrastructure services, including Crossref, Supabase, Render, and ClustrMaps.

### 13 Apr 2026 v2

# A Real-Source Research Workbench for Elite Business Journal Search, Analysis, and Writing Support

@@ Abstract
This paper presents Lishu, a deployable web artifact for searching, monitoring, and interpreting literature from elite business and management journals. The system integrates the UTD-24 and Financial Times 50 (FT50) journal pools and combines Crossref, OpenAlex, Unpaywall, and optional CORE enrichment to support a broader research workflow than article retrieval alone. In the current implementation, users can search across curated journal pools, apply multi-journal filters, preview open full-text excerpts when available, generate citations and exports, inspect topic and affiliation structure, produce review drafts, simulate virtual peer review, and assemble grant-oriented research narratives. Unlike static journal directories or general-purpose academic search engines, the artifact is explicitly scoped to high-status management outlets and is designed to support sensemaking tasks that matter to researchers, doctoral students, and lab managers: identifying recent work, surfacing topical concentration, comparing themes, and converting search output into actionable research material. Architecturally, the system emphasizes source transparency, modularity, and low-cost public deployability through a lightweight Node.js service layer, a multi-page client interface, optional large-language-model enhancement for interpretation and writing support, and a free-tier persistence path through Supabase. The paper contributes both a functioning design artifact and an extensible architectural pattern for journal-pool-specific scholarly discovery and writing support, with implications for digital research infrastructure in information systems and business scholarship.
Keywords: design science, scholarly search, research analytics, business journals, Crossref, OpenAlex, digital research infrastructure
@@ 1 Introduction
This paper introduces Lishu, a web-based research artifact designed to reduce that fragmentation. The artifact does not attempt to replace publisher websites or commercial bibliographic databases. Rather, it provides a focused, domain-specific decision support layer that combines journal-pool curation, real-source metadata retrieval, full-text availability enrichment, topic extraction, export utilities, review-generation support, and lightweight behavioral analytics. In design-science terms, the artifact addresses a class of recurrent scholarly search and sensemaking problems by creating a practical digital object for use in a specific problem domain (Hevner et al., 2004; Gregor and Hevner, 2013). The broader ambition is to show that even a relatively lightweight artifact can offer meaningful infrastructure value when its scope, sourcing logic, and task support are carefully aligned.
@@ 3 Design Objectives
Practical research support. Citation export, favorites, downloadable CSV output, full-text preview, paper interpretation, and structured writing assistance are included because they reduce friction in everyday research tasks.
@@ 4.1 Front-End Structure
    Article Search: multi-source article retrieval, filters, full-text access, hotspot analysis, export, favorites, and citation support.
@@
    Comparative and Daily Views: topic comparison, daily top-paper briefing, and author leaderboard analytics.
@@
    Research Writing Tools: review generation, virtual review, paper reading, grant-proposal drafting, and AutoResearch orchestration.
  6. 6.
    Resources and Analytics: conferences, submission tools, usage analytics, and support materials.
@@ 4.2 Back-End Services
    metadata retrieval, enrichment, and normalization,
@@
Presentation layer HTML, CSS, vanilla JavaScript Multi-page interface for search, journal browsing, writing support, export utilities, and analytics views
Search interaction Browser-side form state and query serialization Stable query submission, multi-journal filtering, shareable URLs, local favorites, and citation actions
@@
Application layer Node.js HTTP server Static asset delivery, API routing, metadata orchestration, writing workflow jobs, and result normalization
Metadata sources Crossref REST API + OpenAlex Live retrieval of paper-level metadata, dates, authors, affiliations, citations, and DOI records
Full-text enrichment Unpaywall + optional CORE Open-access location resolution, PDF availability, and excerpt-level preview support
@@
Presentation Layer Home    Search    Daily Briefing    Topic Comparison    Writing Workflows    Analytics Application Layer Node.js server for routing, Crossref/OpenAlex orchestration, OA enrichment, metadata normalization, query filtering, pagination, citation formatting, job handling, and analytics aggregation Analytics and Workflow Layer Hotspot extraction    distributional analytics    full-text preview    citation export    favorites ranking    optional LLM enhancement Curated Sources UTD-24 metadata FT50 metadata External Metadata Crossref    OpenAlex Unpaywall    CORE Persistence Supabase analytics store or local fallback
Figure 1: System architecture of the Lishu research workbench artifact
@@ 5 Search-to-Insight Workflow
Given a keyword and optional filters, the back end first identifies the journals that belong to the selected pool. It then issues bounded Crossref requests journal by journal, supplements results with OpenAlex records when available, enriches access metadata through Unpaywall and optional CORE retrieval, normalizes the returned data into a common article schema, applies filtering and sorting logic, and computes analytical summaries from the resulting set.
@@
The implemented interface now supports a broader set of researcher-facing tasks than the initial version of the artifact. In addition to journal-pool search and analytics, the system includes daily briefing views, topic-comparison dashboards, review-generation interfaces, paper-reading assistance, proposal drafting, and AutoResearch orchestration. Instead of emphasizing low-fidelity mockups, the report therefore highlights the implemented views that researchers actually interact with during search, synthesis, and writing support. Figures 37 serve not only as screenshots, but as evidence of how the artifact translates architectural intentions into concrete workflows.
@@
Figure 4: Implemented retrieval workspace showing live article search, multi-journal filtering, full-text preview entry points, and research analytics
Figure 5: Implemented comparison and daily-briefing views showing topic timeline analysis and top-paper briefing workflows
Figure 6: Implemented writing-support views showing review generation and grant-proposal composition workflows
Figure 7: Implemented AI-assisted workflow tools showing paper reading and AutoResearch orchestration
Figure 8: Implemented author and analytics views showing author-network exploration and portal-level usage monitoring
Figure 3 illustrates the artifact’s front-door logic: it foregrounds source transparency, journal-pool framing, and navigational shortcuts. Figure 4 shows the core retrieval surface, where live metadata retrieval, filtering, full-text preview, citation export, and hotspot analysis are tightly coupled in a single page. Figure 5 extends the artifact into ongoing field monitoring through topic-comparison and daily-briefing views. Figure 6 shows that the artifact now supports structured downstream writing tasks such as literature review generation and proposal drafting. Figure 7 highlights AI-assisted reading and orchestration workflows, whereas Figure 8 captures both author-level exploration and portal-level usage monitoring. Together, these implemented views demonstrate that the artifact is not only technically functional but also structurally aligned with recurring scholarly work practices.
@@ 6.1 Hotspot Extraction
Hotspot phrases are derived from titles and abstracts of the current result set. Candidate phrases are scored and then grouped into interpretable categories such as AI/Algorithms, Consumer/Market, Operations/Supply Chain, Finance/Accounting, and Strategy/Organization. When no native article keywords are available, the system falls back to rule-based extraction. When an LLM endpoint is configured, hotspot labels and quick takes can be rewritten into more readable summaries. The current interface also surfaces full-text availability signals such as open-access status, PDF availability, source priority, and in-page excerpt preview when an open version can be located.
@@ 7.1 Technology Stack
The artifact uses Node.js for server-side routing and API orchestration, vanilla JavaScript for front-end logic, HTML/CSS for multi-page rendering, ECharts for result-set and analytics visualization, Supabase for free-tier persistence of usage analytics, and Render as a practical public deployment target. The live retrieval layer integrates Crossref, OpenAlex, Unpaywall, and optional CORE enrichment. The implementation deliberately avoids a heavier front-end framework in order to minimize setup friction and hosting complexity. This restraint is important for the artifact’s design logic: a research support system is more likely to be adopted, replicated, and extended when the technical barrier to deployment remains low.
@@ 7.2 Front-End and Interaction Design
The front end is organized as a modular multi-page interface rather than a monolithic single-page application. This choice simplifies routing, page ownership, and public hosting while preserving a coherent user journey across home, journal browsing, live search, topic comparison, daily briefing, review generation, virtual review, paper reading, grant writing, AutoResearch orchestration, and analytics. Browser-side logic manages form state, URL synchronization, multi-journal filtering, favorites persistence, citation actions, and optional user-provided large-language-model configuration. From a software-engineering perspective, this structure improves maintainability by keeping task-specific rendering concerns loosely coupled while still allowing shared interface patterns across pages.
@@ 7.3 Back-End Retrieval and Normalization
The back end mediates between curated journal pools and live metadata retrieval. For each query, the service first resolves the relevant journal set, then issues bounded Crossref requests on a journal-by-journal basis, supplements records with OpenAlex where available, enriches full-text access candidates through Unpaywall and optional CORE lookup, normalizes returned records into a common schema, removes duplicates, applies filters, and computes analytical summaries. This orchestration layer is central to the artifact’s contribution: it operationalizes prestige-based journal curation as a search boundary, thereby aligning retrieval behavior with the evaluative logic commonly used in business schools.
@@ 7.4 Data Provenance and Source Integrity
The artifact is intentionally conservative about data provenance. Journal pools are anchored in documented UTD-24 and FT50 sources, while paper-level content is derived from Crossref and OpenAlex records and exposed through DOI links and source labels. Open-access and full-text hints are derived from Unpaywall and optional CORE enrichment rather than opaque scraping. When article keywords are unavailable from native metadata, the system falls back to explicit heuristic extraction rather than presenting opaque or untraceable enrichment. Optional LLM support is treated as an interpretive and writing enhancement layer rather than a source of authoritative bibliographic truth.
@@ 7.6 Citation and Export Support
Each article object can be exported or cited through several output modes, including BibTeX, APA, MLA, Chicago, and plain text. CSV export and report export are also supported. Beyond search, the current artifact also exposes higher-level writing workflows such as review generation, virtual review simulation, proposal drafting, and AutoResearch-style planning. These functions position the artifact closer to an everyday research workbench than a demonstration-only prototype. In practice, they reduce a common break in scholarly workflow: the transition from discovering relevant papers to incorporating them into notes, bibliographies, analytical dashboards, and downstream writing tasks.
@@ 8 Discussion
The artifact contributes to design science and scholarly infrastructure in three ways. First, it demonstrates a domain-specific scholarly search artifact centered on elite business journals rather than undifferentiated bibliographic breadth. Second, it shows how real-source metadata, open-access enrichment, and lightweight analytics can be combined into a usable search-to-insight workflow without requiring proprietary infrastructure. Third, it extends that workflow into writing-support tasks such as review generation, proposal drafting, and AI-assisted reading while preserving traceable source boundaries. More broadly, the artifact suggests that there is substantial room for middle-layer scholarly systems: systems that do not compete with large bibliographic platforms on scale, but outperform them on domain alignment, transparency, and task focus.
At the same time, the artifact has important limitations. Coverage still depends on the quality and timeliness of Crossref, OpenAlex, and open-access availability services, which vary across journals and articles. The hotspot layer is optimized for interpretability rather than formal topic-model validity. Although the system can preview open-access text excerpts, it does not yet provide comprehensive full-text indexing or citation-network exploration. The higher-level writing modules are useful as research support, but they should be understood as scaffolding and synthesis aids rather than substitutes for scholarly judgment. Finally, the artifact has not yet undergone formal user evaluation with faculty or doctoral students. These limits should be interpreted as research opportunities rather than as disqualifying weaknesses: they point to the next design cycle needed to move from promising artifact to evaluated scholarly infrastructure.
@@ 10 Conclusion
Lishu is a focused research artifact for discovering, searching, and analyzing literature from elite business journals. By integrating UTD-24 and FT50 curation with live multi-source retrieval, open-access enrichment, analytical summaries, optional AI enhancement, citation utilities, and deployable analytics infrastructure, the artifact offers a practical response to the fragmented search workflows often faced by management researchers. The current system also demonstrates how a search artifact can evolve into a broader research workbench that supports reading, comparison, review generation, and proposal writing without abandoning source transparency. Although it remains an early-stage artifact rather than a finished scholarly platform, it already demonstrates a credible and extensible design pattern for business-school-oriented digital research infrastructure. For ICIS-style design and artifact discussions, the promise of the system lies not only in what it currently does, but in the evaluative and infrastructural research program it makes possible.
@@ Acknowledgments
The current public-facing artifact credits Designed by Chuang Zhao (ADM Lab). The implementation also integrates external data and infrastructure services, including Crossref, OpenAlex, Unpaywall, CORE, Supabase, Render, and ClustrMaps.